# SpectrumWorld: Artificial Intelligence Foundation for Spectroscopy

### v3

@@ 4 SpectrumLab
System Overview
Figure 3: Overview of SpectrumLab framework.
AI-ready Datasets and AI-solvable Tasks. SpectrumLab is tightly integrated with SpectrumAnnotator, which is responsible for generating high-quality benchmarks from seed datasets collected from diverse sources. In this workflow, SpectrumAnnotatorcurates a wide range of scientifically rigorous benchmarks from these seeds. SpectrumLab then offers a flexible abstraction for users to define and encapsulate specific AI-solvable tasks based on these curated benchmarks. A core abstraction unique to SpectrumLab is the Benchmark Group. Users can combine multiple benchmark instances or select specific subsets to form a Benchmark Group, creating tailored task definitions within a unified framework. By encapsulating benchmarks as tasks, SpectrumLab streamlines the process of task definition and evaluation, ensuring consistency, scalability, and reproducibility across the platform.
@@
Leaderboards. To ensure transparency and reproducibility, SpectrumLab incorporates a comprehensive public leaderboard system that systematically tracks and compares the performance of a wide range of models across all tasks. The leaderboard provides fine-grained reporting, recording each model’s results on both high-level and detailed tasks. The platform currently supports benchmarking for over 20 MLLMs, including prominent open-source models such as InternVL3 (2025) and proprietary models like GPT-4o (2024), across 14 specific tasks.
Modular Design
@@
Model Signal Perception Semantic Generation Avg. TC QE FE ID GR EP PA PP SE FM MR FP IP DnG Perf. Closed-source MLLMs Claude-3.5-Sonnet 96.36 28.33 76.47 71.43 60.00 77.78 76.32 85.29 82.50 69.23 94.59 20.00 0 0 57.78 Claude-3.7-Sonnet 96.36 38.33 86.27 82.14 71.43 88.89 71.05 88.24 82.28 74.36 89.19 20.00 0 5.26 62.02 Claude-4-Sonnet 96.36 35.00 88.24 92.86 62.22 63.89 60.53 76.47 16.25 43.59 64.86 3.33 0 21.05 45.49 Claude-3.5-Haiku 94.55 31.67 50.98 92.86 66.67 75.00 76.32 76.47 67.50 64.10 81.08 10.00 0 0 52.71 Claude-4-Opus 96.36 33.33 86.27 92.86 73.33 83.33 71.05 85.29 32.50 76.92 86.49 16.67 0 5.26 54.15 GPT-4o 96.36 33.33 68.63 92.86 57.78 77.78 63.16 79.41 78.75 58.97 89.19 10.00 0 0 54.58 GPT-4.1 94.55 28.33 86.27 85.71 53.33 77.78 63.16 79.41 82.50 66.67 91.89 33.33 10.53 0 58.33 GPT-4-Vision 94.55 33.33 72.55 92.86 73.33 72.22 71.05 82.35 73.75 53.85 97.30 23.33 5.00 0 57.51 Grok-2-Vision 94.55 31.67 74.51 89.29 64.44 80.56 73.68 82.35 37.50 66.67 81.08 23.33 0 0 52.21 Qwen-VL-Max 94.55 36.67 90.20 92.86 60.00 80.56 78.95 88.24 32.50 71.79 91.89 43.33 0 5.26 56.64 Doubao-1.5-Vision-Pro 98.18 33.33 78.43 92.86 66.67 83.33 68.42 88.24 67.50 56.41 89.19 6.67 0 0 55.65 Doubao-1.5-Vision-Pro-Thinking 96.36 35.00 78.43 67.86 53.33 80.56 73.68 91.18 68.75 66.67 91.89 66.67 5.00 5.26 61.01 Open-source MLLMs Qwen2.5-VL-32B-Instruct 92.73 26.67 37.25 71.43 57.78 44.44 31.58 61.76 0.00 5.13 45.95 20.00 0 0 30.75 Qwen2.5-VL-72B-Instruct 94.55 38.33 86.27 92.86 42.22 80.56 78.95 88.24 66.25 76.92 91.89 30.00 0 10.53 58.95 InternVL3-78B 96.36 38.33 70.59 71.43 48.49 75.00 81.58 88.24 62.50 69.23 83.78 23.33 0 5.26 55.19 Llama-3.2-11B-Vision-Instruct 34.55 11.67 13.73 25.00 20.00 41.67 15.79 29.41 7.50 5.13 21.62 0 0 0 14.26 Llama-3.2-90B-Vision-Instruct 38.18 10.00 35.29 25.00 17.78 27.78 28.95 20.59 21.25 5.13 43.24 0 0 0 18.78 DeepSeek-VL2 52.73 23.33 29.41 28.57 8.89 27.78 28.95 50.00 15.00 15.38 32.43 10.00 5.00 5.26 22.26 Overall Avg. 86.35 29.81 67.21 75.60 53.21 68.83 61.29 74.51 49.71 52.56 75.98 23.63 1.42 3.51 49.54
@@
Table 3: Accuracies (%, $\uparrow$) of all models on different levels. Task abbreviations (e.g., TC, QE, FE, etc.) are defined in Table 2. best: bold, second best: underlined.
@@
(1) Benchmark Group: SpectrumLab organizes SpectrumBench datasets into hierarchical groups corresponding to different levels of spectroscopic reasoning. This structure supports a broad array of spectroscopic modalities and task types, enabling systematic evaluation across the full spectrum of scientific challenges. The layered grouping not only facilitates comprehensive benchmarking but also allows for rapid and targeted assessment of specialized models on domain-specific spectra and tasks.
(2) Model Integration: SpectrumLab offers a unified and extensible framework for integrating external models. Through standardized APIs and modular adapters, we provide seamless connection to a broad spectrum of model types, from cloud-based services to locally deployed solutions. This enables consistent benchmarking within a single evaluation environment.
@@ 5 Experiment
Benchmark Setup
Evaluation. For signal-, perception-, and semantic-level tasks, SpectrumBench  standardizes them into a multiple-choice question format, with each question having four options. A correct answer is scored as 1, and an incorrect answer is scored as 0. Generation-level tasks usually do not have fixed-form answers. For Molecule-to-Spectrum tasks, the input is a molecule , and the output is a spectrum. For Spectrum-to-Molecule tasks, the input consists of multiple spectral images, and the output is a molecule. We aim to encourage models to generate meaningful reasoning trajectories rather than simply providing a final answer. This approach can help circumvent the issue of data leakage. Therefore, we use an additional MLLM to score the responses following these steps: (1) Model predictions that do not conform to the specified output format for a given question are assigned a score of zero. (2) For predictions meeting the required format, a dedicated scoring model evaluates the model’s output against the answer, assigning a score normalized between 0 and 1. GPT-4o is employed as the scoring model in our experiment. This design standardizes the primary evaluation metric across all tasks in SpectrumBench to accuracy (%).
MLLMs. Leveraging SpectrumLab’s flexible model interface, we integrated 18 leading open- and closed-source MLLMs for our experiments. Further details on benchmarking candidates and a comprehensive cost analysis are provided in Appendix C and G, respectively.
Implementation Details. SwanLab [^2] https://docs.swanlab.cn/en/ was utilized to monitor inference time consumption and evaluation results, with detailed metrics logged and available for review via our anonymous link.
Results & Analysis
@@
(1) Closed-source Models Lead Overall Performance with Claude-3.7-Sonnet Achieving Best Results. Claude-3.7-Sonnet emerges as the top-performing model with an overall average accuracy of 62.02%, securing top-2 scores in 7 out of 14 tasks. The model demonstrates exceptional capabilities across multiple dimensions: it leads in Quality Assessment (QE) with 38.33%, Elemental Compositional Prediction (EP) with 88.89%, and ranks second in several other tasks, including Spectrum Type Classification (TC) and Basic Property Prediction (PP). Closed-source models generally maintain a performance advantage, with the top 5 models by overall average all being proprietary solutions. However, this gap is narrowing, models like Qwen2.5-VL-72B-Instruct(58.95%) and InternVL3-78B (55.19%) are approaching or even surpassing some closed-source counterparts in specific benchmarks.
(2) Reasoning Capabilities Drive Generation Task Performance. Doubao-1.5-Vision-Pro-Thinking demonstrates exceptional performance in generation tasks, achieving 66.67% accuracy in Forward Problems (FP), significantly outperforming the best closed-source model (Qwen-VL-Max at 43.33%). This remarkable 23.34% point advantage highlights the critical role of advanced reasoning capabilities in complex molecule generation tasks. The model also excels in semantic understanding, achieving 68.75% in Molecular Structure Elucidation (SE), 66.67% in Fusing Spectroscopic Modalities (FM), and 91.89% in Multimodal Molecular Reasoning (MR). This superior performance suggests that the “thinking” mode, is essential for tackling sophisticated cross-modal scientific reasoning challenges.
@@
(3) Task Complexity Reveals Model Capabilities and Limitations. Models exhibit strong foundational capabilities in basic tasks, with Signal and Perception tasks showing robust performance across all models. TC achieves an average accuracy of 86.35%, while Impurity Peak Detection (ID) shows an average of 75.60%. However, performance significantly declines in more complex tasks, particularly within the Generation category, which shows an average accuracy of only 23.63%. Within the Generation level, there are notable performance differences: FP achieves an average of 23.63%, significantly outperforming Inverse Problems (IP) at 1.42% and De Novo Generation (DnG) at 3.51%. This suggests that models are more adept at forward prediction tasks (molecule-to-spectrum). QE tasks prove particularly challenging, with an average of 29.81% across all models, and many models scoring 0% in IP and DnG tasks. This performance pattern reveals a clear hierarchy: models excel at basic pattern recognition and signal processing but struggle with advanced reasoning, creative generation, and complex cross-modal synthesis tasks that require deeper scientific understanding.
(4) Parameter Scaling and Architecture Optimization Show Clear Benefits. The Qwen2.5-VL series demonstrates the significant impact of model scaling, with Qwen2.5-VL-72B-Instruct (58.95%) achieving nearly double the performance of its 32B counterpart (30.75%). This dramatic improvement highlights the importance of parameter count and architectural sophistication for complex multimodal reasoning tasks. The scaling effect is particularly evident in semantic level, where the larger model shows substantial gains in Molecular Structure Elucidation (66.25% vs 0.00%) and Fusing Spectroscopic Modalities (76.92% vs 5.13%). These results suggest that architectural optimization and training improvements hold significant promise for advancing spectroscopic AI capabilities.
@@ 6 Conclusion
In this work, we present two key contributions to advance machine learning in spectroscopy: SpectrumBench and SpectrumLab. SpectrumBench is a comprehensive, extensible benchmark suite covering over 10 spectrum modalities and 14 tasks, grounded in real-world chemical practices, enabling rigorous and reproducible evaluation across hierarchical taxonomy (signal, perception, semantic, generation). SpectrumLab is a unified, modular platform for dataset management, annotation, evaluation, and public leaderboards, offering a robust Python ecosystem with standardized interfaces that significantly lower the barrier for developing and deploying advanced models. Together, SpectrumBench and SpectrumLab set a new standard for spectroscopic machine learning, fostering systematic comparison, reproducibility, and innovation, and catalyzing future research for more powerful and interpretable models.
Appendix A SpectrumBench Detailed Information
Signal Level

### v4

@@ 4 SpectrumLab
4.1 System Overview
AI-ready Datasets and AI-solvable Tasks. SpectrumLab is tightly integrated with SpectrumAnnotator, which is responsible for generating high-quality benchmarks from seed datasets collected from diverse sources. In this workflow, SpectrumAnnotator curates a wide range of scientifically rigorous benchmarks from these seeds. SpectrumLab then offers a flexible abstraction for users to define and encapsulate specific AI-solvable tasks based on these curated benchmarks. A core abstraction unique to SpectrumLab is the Benchmark Group. Users can combine multiple benchmark instances or select specific subsets to form a Benchmark Group, creating tailored task definitions within a unified framework. By encapsulating benchmarks as tasks, SpectrumLab streamlines the process of task definition and evaluation.
@@ 4.1 System Overview
Leaderboards. To ensure transparency and reproducibility, SpectrumLab incorporates a comprehensive public leaderboard system that systematically tracks and compares the performance of a wide range of models across all tasks. The leaderboard provides fine-grained reporting, recording each model’s results on both high-level and detailed tasks. The platform currently supports benchmarking for over 20 MLLMs, including prominent open-source models such as InternVL3 (Zhu et al., 2025) and proprietary models like GPT-4o (OpenAI, 2024), across 14 specific tasks.
4.2 Modular Design
@@
Model Signal Perception Semantic Generation Avg. TC QE FE ID GR EP PA PP SE FM MR FP IP DnG Perf Closed-source MLLMs Claude-3.5-Sonnet 96.36 28.33 76.47 71.43 60.00 77.78 76.32 85.29 82.50 69.23 94.59 20.00 0 0 59.88 Claude-3.7-Sonnet 96.36 38.33 86.27 82.14 71.43 88.89 71.05 88.24 82.28 74.36 89.19 20.00 0 5.26 63.84 Claude-4-Sonnet 96.36 35.00 88.24 92.86 62.22 63.89 60.53 76.47 16.25 43.59 64.86 3.33 0 21.05 51.76 Claude-3.5-Haiku 94.55 31.67 50.98 92.86 66.67 75.00 76.32 76.47 67.50 64.10 81.08 10.00 0 0 56.23 Claude-4-Opus 96.36 33.33 86.27 92.86 73.33 83.33 71.05 85.29 32.50 76.92 86.49 16.67 0 5.26 59.98 GPT-4o 96.36 33.33 68.63 92.86 59.88 77.78 63.16 79.41 78.75 58.97 89.19 10.00 0 0 57.74 GPT-4.1 94.55 28.33 86.27 85.71 53.33 77.78 63.16 79.41 82.50 66.67 91.89 33.33 10.53 0 60.96 GPT-4-Vision 94.55 33.33 72.55 92.86 73.33 72.22 71.05 82.35 73.75 53.85 97.30 23.33 5.00 0 60.39 Gemini-2.5-pro 96.36 35.00 90.20 67.86 75.56 86.11 65.79 79.41 68.75 84.62 97.30 50.00 5.00 47.37 67.81 Grok-2-Vision 94.55 31.67 74.51 89.29 64.44 80.56 73.68 82.35 37.50 66.67 81.08 23.33 0 0 57.12 Qwen-VL-Max 94.55 36.67 90.20 92.86 60.00 80.56 78.95 88.24 32.50 71.79 91.89 43.33 0 5.26 61.91 Doubao-1.5-Vision-Pro 98.18 33.33 78.43 92.86 66.67 83.33 68.42 88.24 67.50 56.41 89.19 6.67 0 0 59.23 Doubao-1.5-Vision-Pro-Thinking 96.36 35.00 78.43 67.86 53.33 80.56 73.68 91.18 68.75 66.67 91.89 66.67 5.00 5.26 62.90 Open-source MLLMs Qwen2.5-VL-32B-Instruct 92.73 26.67 37.25 71.43 57.78 44.44 31.58 61.76 0.00 5.13 45.95 20.00 0 0 35.34 Qwen2.5-VL-72B-Instruct 94.55 38.33 86.27 92.86 42.22 80.56 78.95 88.24 66.25 76.92 91.89 30.00 0 10.53 62.68 InternVL3-78B 96.36 38.33 70.59 71.43 48.49 75.00 81.58 88.24 62.50 69.23 83.78 23.33 0 5.26 58.15 InternVL3.5-241B 98.18 33.33 90.20 92.86 66.67 77.78 71.05 85.29 86.25 61.54 100.00 33.33 10.00 10.53 65.50 Llama-3.2-11B-Vision-Instruct 34.55 11.67 13.73 25.00 20.00 41.67 15.79 29.41 7.50 5.13 21.62 0 0 0 16.15 Llama-3.2-90B-Vision-Instruct 38.18 10.00 35.29 25.00 17.78 27.78 28.95 20.59 21.25 5.13 43.24 0 0 0 19.51 DeepSeek-VL2 52.73 23.33 29.41 28.57 8.89 27.78 28.95 50.00 15.00 15.38 32.43 10.00 5.00 5.26 23.77 GLM-4.5V 100.00 28.33 70.59 92.86 73.33 83.33 71.05 91.18 63.75 69.23 83.78 0 0 10.53 59.85 InternS1-nothink 98.18 36.67 72.55 89.29 51.11 72.22 73.68 79.41 86.25 66.67 94.59 13.33 0 0 59.57 InternS1-think 98.18 25.00 80.39 89.29 64.44 88.89 73.68 91.18 90.00 56.41 91.89 10.00 40.00 15.79 65.37 Overall Avg. 89.09 30.65 70.16 77.95 56.13 71.62 63.84 76.85 56.08 55.85 79.79 20.29 3.50 6.41 54.15
@@
Table 3: Accuracies (%, $\uparrow$) of all models on different levels. Task abbreviations (e.g., TC, QE, FE, etc.) are defined in Table 2. best: bold, second best: underlined. The second last column calculates the arithmetic mean and the last column calculates the true weighted mean of each row.
@@
(1) Benchmark Group: SpectrumLab organizes SpectrumBench into hierarchical groups corresponding to different levels of spectroscopic reasoning. This structure enables systematic evaluation across various tasks and spectroscopic modalities, while also supporting rapid assessment of specialized models on domain-specific spectra. spectra and tasks.
(2) Model Integration: SpectrumLab offers a unified, extensible framework for integrating external models. Using standardized APIs and modular adapters, it connects seamlessly to a wide range of model types, from cloud-based services to locally deployed solutions, enabling consistent benchmarking within a single evaluation environment.
@@ 5 Experiment
5.1 Benchmark Setup
For signal-, perception-, and semantic-level tasks, SpectrumBench standardizes them into a multiple-choice question format, with each question having four options. A correct answer is scored as 1, and an incorrect answer is scored as 0. Generation-level tasks usually do not have fixed-form answers. For Molecule-to-Spectrum tasks, the input is a molecule, and the output is a spectrum. For Spectrum-to-Molecule tasks, the input consists of multiple spectral images, and the output is a molecule. We aim to encourage models to generate meaningful reasoning trajectories rather than simply providing a final answer. This approach can help circumvent the issue of data leakage. Therefore, we use an additional MLLM to score the responses following these steps: (1) Model predictions that do not conform to the specified output format for a given question are assigned a score of zero. (2) For predictions meeting the required format, a dedicated scoring model evaluates the model’s output against the answer, assigning a score normalized between 0 and 1. GPT-4o is employed as the scoring model in our experiment. This design standardizes the primary evaluation metric across all tasks in SpectrumBench to accuracy (%). Leveraging SpectrumLab’s flexible model interface, we integrated 23 leading open- and closed-source MLLMs for our experiments. Further details on benchmarking candidates and cost analysis are provided in Appendix 14 and H, respectively.
5.2 Main Findings
@@
(1) Task complexity reveals model capabilities and limitations. Models exhibit strong foundational capabilities in basic tasks, with Signal and Perception tasks showing robust performance across all models. Spectrum Type Classification(TC) achieves an average accuracy of 89.09%, while Impurity Peak Detection (ID) shows an average of 77.95%. However, performance significantly declines in more complex tasks, particularly within the Generation category, which shows an average accuracy of only 6.41%. Within the Generation level, there are notable performance differences: FP achieves an average of 20.29%, significantly outperforming Inverse Problems (IP) at 3.50% and De Novo Generation (DnG) at 6.41%. This suggests that models are more adept at forward prediction tasks (molecule-to-spectrum) than inverse problems. QE tasks prove particularly challenging, with an average of 30.65% across all models, and many models scoring 0% in IP and DnG tasks. This performance pattern reveals a clear hierarchy: models excel at basic pattern recognition and signal processing but struggle with advanced reasoning, creative generation, and complex cross-modal synthesis tasks that require deeper scientific understanding.
(2) Closed-source models lead overall performance with gemini-2.5-pro achieving best results. Gemini-2.5-pro emerges as the top-performing model with an overall average accuracy of 67.81%, securing top-2 scores in 6 out of 14 tasks. The model demonstrates exceptional capabilities across multiple dimensions: it leads in Functional Group Recognition (GR) with 75.56%, and ranks second in several other tasks, including Elemental Compositional Prediction(EP) and Forward Problems(FP). Closed-source models generally maintain a performance advantage. However, this gap is narrowing, models like InternVL3.5-241B(65.50%) and InternS1-think (65.37%) are approaching or even surpassing some closed-source counterparts.
(3) Reasoning capabilities drive generation task performance. Doubao-1.5-Vision-Pro-Thinking demonstrates exceptional performance in generation tasks, achieving 66.67% accuracy in Forward Problems (FP), significantly outperforming the second-best closed-source model (Gemini-2.5-pro at 50.00%). This remarkable 16.67% point advantage highlights the critical role of advanced reasoning capabilities in complex molecule generation tasks. InternS1-think also outperforms InternS1 (65.37% vs. 59.57%). This superior performance suggests that the “thinking” mode is essential for tackling sophisticated cross-modal scientific reasoning challenges.
5.3 Error Analysis
Figure 5: Error types and distributions in SpectrumBench.
Error Analysis.
For our error analysis, we group the 14 tasks in SpectrumBench into five families: Basic recognition (elemental composition, functional-group recognition, spectrum-type classification, spectrum-quality assessment), Peak analysis (peak assignment, impurity-peak detection, basic feature extraction), Property prediction (basic property prediction, molecular structure elucidation), Complex reasoning (multimodal molecular reasoning, fusing spectroscopic modalities, forward/inverse problems), and Generation (de novo spectrum generation).
Family-wise minima. On Basic recognition, Gemini-2.5-Pro attains the lowest error (29.1%), indicating comparatively stronger grounding in spectrum type/quality and functional-group cues. For Peak analysis, Qwen2.5-VL-72B achieves the lowest error (14.5%), suggesting effective handling of isotopic/fragment patterns and impurity peaks. Within Property prediction, Intern-S1 yields the best result (10.5%), followed by InternVL-3.5 (13.2%); both exhibit more reliable mapping from spectral evidence to molecular properties/structures. The Complex reasoning slice is the sub-most challenging: although Gemini-2.5-Pro leads with 30.7% error, the majority of models exceed $50\%$ in this family, underscoring difficulties with long-horizon, cross-modal deduction. For Generation, Gemini-2.5-Pro again performs best (52.6% error), while several models approach failure on nearly all instances (errors near $100\%$).
[…]
@@ 6 Conclusion
In this work, we have presented two key contributions to advance machine learning in spectroscopy: SpectrumBench and SpectrumLab. SpectrumBench is a comprehensive, extensible benchmark suite covering over 10 spectrum modalities and 14 tasks, grounded in real-world chemical practices, enabling rigorous and reproducible evaluation across hierarchical taxonomy (signal, perception, semantic, generation). SpectrumLab is a unified, modular platform for dataset management, annotation, evaluation, and public leaderboards, offering a robust Python ecosystem with standardized interfaces that significantly lower the barrier for developing and deploying advanced models. Together, SpectrumBench and SpectrumLab set a new standard for spectroscopic machine learning, fostering systematic comparison, reproducibility, and innovation, and catalyzing future research for more powerful and interpretable models.
Appendix A Ablation Study
Temperature
Temp Qwen2.5 VL-32B Qwen2.5 VL-72B InternVL3-78B Signal Perception Semantic Generation Signal Perception Semantic Generation Signal Perception Semantic Generation 1.0 57.02 48.89 17.03 6.67 78.00 72.49 78.35 13.51 69.18 73.33 71.84 9.53 0.5 63.40 68.63 29.49 36.23 66.49 65.36 58.97 17.39 66.49 65.36 58.97 17.39 0.0 63.92 70.59 28.85 37.68 63.92 60.78 66.03 17.39 63.92 60.78 66.03 17.39
Table 4: Performance of three models under varying temperature settings.(Top-$p$ fixed to 1)
Table 4 presents the impact of varying temperature settings on three models: Qwen2.5-VL-32B, Qwen2.5-VL-72B, and InternVL3-78B.
For Qwen2.5-VL-32B, lower temperatures ($T=0.5$ and $T=0$) yield substantial improvements over $T=1.0$, particularly on the Perception, Semantic, and Generation levels. A similar trend is observed for InternVL3-78B, where deterministic decoding ($T=0$ or $T=0.5$) leads to a more balanced performance profile compared to the stochastic setting. In contrast, Qwen2.5-VL-72B behaves differently: while it achieves the highest Signal and Semantic scores at $T=1.0$, its Generation accuracy remains relatively low across all settings.
These observations indicate that smaller models tend to benefit from reduced sampling variability, as lower temperatures enhance stability and reliability. Conversely, larger models may require higher temperatures to fully exploit their expressive capacity, though this comes at the cost of weaker generative consistency.
Top-$p$ Qwen2.5 VL-32B Qwen2.5 VL-72B InternVL3-78B Signal Perception Semantic Generation Signal Perception Semantic Generation Signal Perception Semantic Generation 1.0 57.02 48.89 17.03 6.67 78.00 72.49 78.35 13.51 69.18 73.33 71.84 9.53 0.5 63.92 69.93 29.49 39.13 62.89 62.09 65.38 13.04 70.62 68.63 72.44 5.80 0.1 63.40 66.67 30.77 44.93 66.49 54.25 64.74 23.19 69.07 69.93 72.44 11.59
Table 5: Performance of three models under varying Top-$p$ settings.(temperature fixed to 1).
Top-$p$
We further investigate the role of nucleus sampling while fixing the temperature to 1.0. The results in Table 5 show heterogeneous effects across models.
For Qwen2.5-VL-32B, reducing Top-$p$ from 1.0 to 0.1 consistently improves Semantic and Generation scores, suggesting that constraining the sampling space mitigates low-quality outputs and enhances reliability. By contrast, Qwen2.5-VL-72B attains its best Signal and Semantic results at $p=1.0$, but its Generation score is substantially reduced. Interestingly, setting $p=0.1$ recovers part of this loss, implying a trade-off between precision and diversity.
For InternVL3-78B, performance remains comparatively stable across Top-$p$ values, with minor fluctuations in Generation accuracy. This stability suggests that larger-scale models are less sensitive to sampling truncation, reflecting stronger intrinsic consistency.
Scoring Model
The automatic evaluation at the Generation level uses a standardized prompt to guide the scoring models. This prompt, shown in the box below, instructs the evaluator to rate model answers on a scale from 0 to 1 based on specific rules.
[…]
Scoring Model FP IP DnG
GPT-4o 23.33 0.00 5.26
InternVL3-78B 23.33 5.26 5.00
Table 6: Comparison of different scoring models (GPT-4o vs. InternVL3-78B) on Generation Level evaluation. FP, IP, and DnG represent different sub-categories.
Since the Generation level relies on automatic evaluation, we further compare different scoring models (Table 6). Both GPT-4o and InternVL3-78B yield consistent FP scores (23.33%). However, discrepancies arise in the IP dimension, where GPT-4o assigns zero while InternVL3-78B detects additional errors (5.26%). For DnG, the results are nearly identical (5.26% vs. 5.00%).
These findings suggest that overall evaluation trends are robust across scoring models, but fine-grained categories may be influenced by the evaluator’s internal biases. Consequently, careful selection of the scoring model is essential to ensure fairness and reliability in automatic evaluation pipelines.
Appendix B SpectrumBench Detailed Information
B.1 Signal Level